OpenStudy (anonymous):

the ares of a circle with radius r is given by the formula A=pie*r^2. if a circle has area 225pie sq in. what the radius

OpenStudy (anonymous):

225=pie*r^2 225/pie = r^2 15/sqrt(pie) = r r = 8.4628

OpenStudy (anonymous):

what is 8.4 in a fraction

OpenStudy (anonymous):

\[A=\pi*r^2\] \[225\pi=\pi*r^2\] \[r^2=\frac{225\cancel \pi}{\cancel \pi}\] \[r=\sqrt{225}\] \[r=15\]